Report for July 2016

Report Date
July 2016

Report compiled by Tim Gray, IS Apps Programme Managers for Alex Carter  (Programme Owner)

Section One: Programme Commentary, RAG and Resources Update

Section Two: Projects Not Started

Section Three: Projects In Progress

Section Four: Closed or Withdrawn Projects 

Section Five: Actions from Monthly Meeting

Section One: Programme Commentary, RAG and Resources Update

In a busy month for the programme, the Wiki upgrade project successfully completed the initial part of the upgrade work, upgrading the DB to Oracle 12 and moving the application to new infrastructure.

The only RED project COM022 is undergoing an options appraisal to ascertain the best way forward. Testing and resource availability are the main challenges here.

COM023 has been closed as planned. The BI Architecture project is on schedule to close in August.  

RAG Status and Forward Look

Resource Summary

 Section Two: Projects Not Started

Project InfoProject Manager's Commentary

 

All this year's projects are now in flight.

 

Section Three: Projects in Progress

Project InfoProject Manager's Commentary

COM022 Academic Timetable Integration into Office 365 Calendar  

Discretionary

Milestones 

02-Jun-2016Acceprtance Sign off-Go/no Go
06-Jun-2016Deploy View changes to Live (new)
09-Jun-2016Switch LIVE completed
27-Jun-2016Deployment Sign Off review (DSOR)
13-Jul-2016Close

 

RAG Status

RED

Report by Chris Copner for July 2016

Achievements this period

  • Reviewed outcomes of volume testing

Issues

  • We have not achieved a successful throughput of the test data and there is a significant gap between actual and expected throughput.  The stop/start nature of the throughput has also impacted our ability to measure its speed and performance which is another key test of viability.  The deployment of the software into a Live environment is dependent on running a successful test and we are, at present, unable to provide the acceptance required to proceed.  The team has identified that a second round of testing will be required but this can only commence once we have clarity about the root cause of the issue with the data throughput and an appropriate fix implemented.  The potential reasons for issues that the supplier has provided have been explored at UoE but both the SQL database is working fine, and the views we created have tested successfully.  To meet with sponsor to decide whether to go for a re-test with supplier onsite.

  • Acceptance and Deployment milestones have been passed; and will need to be revised

Next Steps

  • Re-work as required, organise supplier onsite if sponsor wishes to pursue
  • Document Operational Level Agreement (OLA)

 

COM025

BI Architechture

Discretionary

Milestones

19-Aug-2016'To Be' Model Sign Off
26-Aug-2016Closure
26-Aug-2016Final Report

RAG

GREEN

Report by Craig Middlemass for July 2016

Achievements in the last period

  • The To-Be model and roadmap have been circulated for comment
  • Some feedback has been recieved on the roadmap and Service Level Agreement
  • Several handover discussions have taken place with the COM029 project team which will deliver the core of a Enterprise Data Warehouse

Issues

  • For info - A large portion of July the project team were occupied on higher priority work and annual leave - the expectation is that the remaining sign off will occur quickly

Next steps

  • Gather feedback and sign off the SLA
  • Gather feedback and sign off the Roadmap
  • Gather feedback and sign off the To-be model
  • Estimate the Service Requirements
  • Integrate the work from the IS BI Working Group into templates for Annual Planning, Business Requirement, Analysis and Design work

 

 

COM026

Wiki Upgrade

Discretionary

Milestones

01-Jul-2016Integration reviewBuild
22-Jul-2016Acceptance reviewAccept
25-Jul-2016DeliveryDeliver
15-Aug-2016Deployment reviewDeliver
31-Aug-2016ClosureClose

 

RAG Status

AMBER

Report by Tim Gray for July 2016

Achievements in Last Period

  • TEST upgrade completed
  • Load testing completed
  • UAT completed
  • DB upgrade and move to new hardware successfully completed on 30-31/07/16

Issues

  • No significant issues to report,

Next Steps

  • Upgrade new LIVE environment to new version
  • Move to closure

COM027

Enterprise Data services Roll Out

COM027

WIKI Upgrade

Discretionary

Milestones:

   13-Jul-16                       Planning Review (see PICCLE).

RAG

GREEN

Report by Iain Moore-McLaughlin for July 2016

Achievements in Last Period

  • Hand over of project management responsibilities complete.
  • Project approach agreed; project to focus on Events Booking initially subject to agreement with Student Systems.
  • Regular weekly checkpoint meetings scheduled with Project Sponsor.

Issues

  • Project Brief has not yet been completed and signed-off; 25% of the planned budget has been utilised.

Next Steps 

  • Project Brief to be completed and signed-off.
  • Project estimates and project plan to be validated and updated as necessary.
  • Milestones to be created and agreed.
  • Project resources to be requested.
  • Work on Technical Architecture Document to be deferred until after Project Brief Review / Planning meeting on 11th August. 
  • Project Brief Review / Planning meeting scheduled 11th August - the outputs from this meeting will inform completion of the Project Brief, validation of project estimates, project plan and creation of milestones
  • Planning meeting with Student Systems rescheduled to 18th August.

Section Four: Closed or Withdrawn Projects

Project InfoProject Managers Commentary

 

COM019

IDM MIgration to Oracle 11G

Compliance

Report by Colin Forrest

This project was closed on 21/08/15 (WIS Notes 21/08/15).

Please see the Completion Report for further information.

COM014

Update and Rationalise Windows Web Hosting

Discretionary

Report by David Watters

This project was closed on 21/08/15 (WIS Notes 21/08/15)

Please see the Completion Report for further information

COM017

Wiki Service Improvement

Discretionary

Report by David Watters

This project was closed on 02/10/15 (WIS Notes 02/10/15)

Please see the Completion Report for further information

COM020

SharePoint Digital Space for IS Consolidation

Discretionary

Report by Colin Watt

This project was closed on 30/10/15 (WIS Notes 30/10/15)

Please see the Completion Report for further information

COM018

VRS Improvements

Discretionary

Report by Chris Konzak

This project was closed on 11/12/2015 (WIS Notes 11/12/2015 )

Please see the Completion Report for further information

AP56-002 Upgrade onsite Exchange Mail Relay Servers

Discretionary

Project withdrawn on 27/11/15 see PRICCL 23 and WIS Notes 27/11/15

COM016

Committee Intranet Migration

Discretionary

Report by Tim Gray

This project was closed on 01/04/2016 (WIS Notes 01/04/2016)

Please see the Completion Report for further information

COM021

IDM Continuous Improvements

Discretionary

Report by Karen Stirling

This project was closed on 15/04/2016 (WIS Notes 15/04/2016)

Please see the Completion Report for further information

COM015

BI Additional Tools –Dashboards/ReportWriter

Discretionary

Report by Tim Gray

This project was closed on 29/04/2016 (WIS Notes 29/04/2016)

Please see the Completion Report for further information

COM023

Enhancements to Visitor Registration System (including CF Upgrade)

Discretionary

Report by Sue Woodger

This project was closed on 10/06/2016  (WIS Notes 29/07/16)

Please see the Completion Report for further information

 

An issue with functional accounts is being taken forward under RDS

COM024

COM Programme CF updates

Discretionary

Report by Sue Woodger

This project was closed on 10/06/2016 (WIS Notes 10/06/16)

Please see the Completion Report for further information

Section Five: Actions/Notes from monthly meeting held on 11/04/15.

 

ProjectAction / NotesResponsibilityStatus
COM023Follow up with PM about functional ageingTGDone 
New EDW Project (COM028)Submit TR code request to WISTGDone - 12.08.16
Forward LookFollowing changes:
  • Confirm budget for COM028 with directors office
  • Move Wiki upgrade to start in Jan
  • BI suite move to Sept - done
  • Add Serengeti upgrade
TG
  • Open
  • Done
  • Done
  • Done